OpenStudy (itsstribe):

Two functions are given : f(x) = -4x+1 g(x)= -4 + 1/2 How does the graph of f compare with the graph of g? A. The graph of f is less steep than the graph of g. B. The graph of f has the same y-intercept as the graph of g. C. The graph of f is paralle to the graph of g. D. The graph of f is steeper than the graph of g.

OpenStudy (itsstribe):

Please help fan and medal.. Very confused with this question!

OpenStudy (welshfella):

there is no x in the function g typo somewhere

OpenStudy (welshfella):

should it be g(x) = -4x + 1/2 ?

OpenStudy (welshfella):

check your original source

OpenStudy (itsstribe):

Ok let me check

OpenStudy (itsstribe):

oh yes it's g(x) = -4x + 1/2

OpenStudy (welshfella):

ok Now functions like this containing x have graphs which are straight lines and theie general form can be written y = mx + b where m = slope of the line and b is the y intercept. Lines of the same slope will be parallel.

OpenStudy (welshfella):

so comparing y = -4x + 1 and y = -4x + 1/2 which of the 4 choices is correct?

OpenStudy (itsstribe):

Hmm..

OpenStudy (itsstribe):

I seriously don't know how to do y=mx+b with those equations, Or to answer the questions

OpenStudy (itsstribe):

Would it be B. The graph of f has the same y intercept as the graph of g?

OpenStudy (welshfella):

the m simply means 'any number' and the same with the b m could be -5, 7 0.12 etc y = mx + b is just a short way of writing 'y equals some number times x + some number' but the m is the values of the slope of the line and b = y intercept

OpenStudy (welshfella):